500 L of a gas at 300 K and 700 Torr would occupy what volume at STP.
1
Expert's answer
2020-04-13T02:26:00-0400
  • According to the conditions of the question we should make up two equations of Mendeleev-Clapeyron, which characterize the initial and final state for this gas:

"p(1) • V(1) = \nu • R • T(1)"

"p(2)\u2022V(2)=\u03bd\u2022R\u2022T(2)"

  • Further, by dividing the first equation by the second, we obtain ("\nu" and R are constants):

"(p(1) • V(1))/ (p(2) • V(2)) = T(1)/T(2) \to"

"V(2) = (p(1) \u2022 V(1) \u2022 T(2)) \/ (p(2) \u2022 T(1))"

  • Considering that: 700 Torr = 93326 Pa, STP : p = 105 Pa, T = 273 K, Pa=N/m2, L = m3, we will get:

V(2) = (93326(N/m2) • 5•103(m3) • 273(K)) / (105(N/m2) • 300(K)) = (127389990•103) / (300•105) = 424633.3•10-2 = 4246.3 m3 = 4246.3 L

Answer: V(2) = 4246.3 L